OpenStudy (anonymous):

f(x)= 2x+1, g(x)= X^2 - 1 what is (fof)(x) and (gof)(x)?

OpenStudy (anonymous):

f composed of f = 2(2x+1) +1 = 4x+3 g composed of f = (2x+1)^2 -1 = 4x^2+4x

OpenStudy (anonymous):

Thanks QED!

OpenStudy (anonymous):

What is gog(x)?

OpenStudy (anonymous):

My answer is x^2-2x i think it was wrong